40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 23 21:53:06 2012, Jenne, Update, Green Locking, Xgreen still wouldn't lock 
    Reply  Thu Aug 23 22:54:32 2012, Jenne, Update, Green Locking, Xgreen still wouldn't lock Xarm_Green_ErrorReflSignals_23Aug2012_LowRes.png
Message ID: 7266     Entry time: Thu Aug 23 22:54:32 2012     In reply to: 7261
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: Xgreen still wouldn't lock 

Quote:

[Jenne, Jamie]

We took a look at the Xend green, and we weren't able to make it lock.  We improved the alignment a little bit, and when we looked at the error signal, it looked nice and PDH-y, but for whatever reason, the cavity won't catch lock.

While aligning the green to the arm, Jamie noticed that the reflection from the intracavity power (not the prompt reflection) was not overlapping with the input beam or prompt reflection.  This means that the cavity axis and the input green beam were not co-linear.  I adjusted the BS and ITMX to get the IR transmitted beam (which had been near clipping on the top edge of the first (2 inch) optic it sees out of the vacuum) back near the input green beam spot on the combining beam splitter.  Then we continued tweaking the green alignment until we saw nice TEM00 flashes in the cavity.  The SNR of the error signal increased significantly after this work, since the cavity buildup was much higher.  But alas, still no lock. 

 I tweaked the alignment of ITMX and ETMX a teeny bit to get the TEM00 flashes back (the work in the previous elog was pre-dinner, so it had been a few hours), then took a screenshot of the error signal and refl dc power on the photodiode for the green xend setup.

The error signal is certainly noisy, although I think when Jamie and I were looking at it earlier this evening, the SNR was a little better.

I need to look at the modulation depth, to see if it's correct, ... maybe lock the Xarm on IR and scan the green laser PZT to check the sideband heights.

I should also check to make sure that the PD is powered, and the gain is high enough (currently the PD gain is set to 20dB).  Earlier today, when I set the gain to 30dB, Jamie said that it was saturating, so I put it back down to the 20dB where we found it.

Still no lock of the green though :(

 

Edit: realized I was bad and didn't label the traces on the plot:  green is refl dc power, blue is demodulated error signal.

Attachment 1: Xarm_Green_ErrorReflSignals_23Aug2012_LowRes.png  96 kB  Uploaded Fri Aug 24 00:23:49 2012  | Hide | Hide all
Xarm_Green_ErrorReflSignals_23Aug2012_LowRes.png
ELOG V3.1.3-